Let T: P2 --> R2 be the linear transformation such that T(x+1)=(1,1), T(x2)=(1,0) and T(x-1)=(0, 1). Find T(2+x+x2).
Answer #1

Sola consider the standard basis B = {1, 2, x²} of P2: Then the coordinate nector Welt + علوم 0 [x+1] = 1 L1 B 0 - 1 2 [x-1361 0 2 By then 0 - 0 0 0 - - 12 0 RTR + R₂ 00 3/2 0 0 0 0 - 12 1 0 0 Thus!! یعنیلی 2 + x + x² = 3 (2+1) + 142?) + ++) (r=1) th